OpenStudy (anonymous):

prove that, for all nonzero a, and b (a * b)^x = a^x * b^xprove that, for all nonzero a, and b (a * b)^x = a^x * b^x@Mathematics

OpenStudy (anonymous):

(ab)^x = e^(xln(ab)) =e^(xln(a)+xlnb) =e^(xln(a))*e^(xln(b)) = a^x * b^x

OpenStudy (anonymous):

(ab)^x = (ab)(ab)...(ab) , rearrange the a's, and b's

OpenStudy (anonymous):

that proof works.

OpenStudy (anonymous):

(ab)^x = (ab)(ab)(ab)...= (aaa...)(bbb...)

OpenStudy (anonymous):

but this only works for positive integer x, so xavier's proof is better

OpenStudy (anonymous):

but using his proof would mean that I have to know what ln means.

OpenStudy (anonymous):

ln is the inverse of e^x

OpenStudy (anonymous):

Yeah. Some properties of logs are used and the fact that e^(a+b)=e^a*e^b

OpenStudy (anonymous):

ok, well then , how can you show that (ab)^(1/3) = (a)^1/3 * b^(1/3)

OpenStudy (anonymous):

xavier, youre proof also works for irrational numbers like pi

OpenStudy (anonymous):

:):):):):):)

OpenStudy (anonymous):

i love pi

OpenStudy (anonymous):

It assumes a lot though. Pecan pi :)

OpenStudy (anonymous):

it assumes things like least upper bound property , but thats fine because real numbers have completeness property

OpenStudy (anonymous):

agdg, are you doing this for a class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we define number as the equivalence class of sets which are 1-1 correspondance with it

OpenStudy (zarkon):

a=-1,b=-1,x=1/2

OpenStudy (anonymous):

\[(ab)^x = ((ab)^1)^x = (a^1b^1)^x = a^x * b^x\]meh my proof is lame... but I'm getting there :-(

OpenStudy (anonymous):

that doesnt follow

OpenStudy (zarkon):

it is hard to prove something that isn't true...you need to put more restrictions on a and b. see my example

OpenStudy (anonymous):

right

OpenStudy (anonymous):

In step 2 to 3 of your proof you used what you are trying to prove

OpenStudy (anonymous):

(ab)^x = e^(xln(ab)) =e^(xln(a)+xlnb) =e^(xln(a))*e^(xln(b)) = a^x * b^x, so a>0 , b>0 here, otherwise it does not work

OpenStudy (anonymous):

because ln a is not define for a <=0

OpenStudy (anonymous):

I can already prove that \[(ab)^{-1} = a^{-1}b^{-1}\]

OpenStudy (anonymous):

yeah thats easy

OpenStudy (anonymous):

i have an interesting question what is -8^(2/3)

OpenStudy (anonymous):

it is -4

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

order of operations i think

OpenStudy (anonymous):

I learned that exponentiation takes precedence over unary minus the hard way :(

OpenStudy (anonymous):

heres another zinger

OpenStudy (anonymous):

(-8)^(4/2)

OpenStudy (anonymous):

64

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

if we do it as [(-8)^(1/2)]^4 that is not defined

OpenStudy (anonymous):

but ((-8)^4)^(1/2) is defined

OpenStudy (anonymous):

if you allow complex numbers it comes out the same though.

OpenStudy (anonymous):

Do parenthesis first?

OpenStudy (anonymous):

oh nice

OpenStudy (anonymous):

:( won't anyone prove for me that (ab)^x = a^x * b^x ? :(

OpenStudy (anonymous):

\[[(-8)^{\frac{1}{2}}]^4=[i\sqrt{8}]^4=i^4\cdot 64=64\]

OpenStudy (anonymous):

I already did

OpenStudy (anonymous):

but you said it won't work for x in (-1, 0) U (0, 1)

OpenStudy (anonymous):

(ab)^x = (ab)(ab)...(ab) x times rearrange by commutativity of numbers (aa....a)(bb...b) where each of them are x times a^x*b^x

OpenStudy (anonymous):

lets just change my question to include for all positive reals x, use Xavier's proof, and call it a day?

OpenStudy (anonymous):

actually it does, for instance (ab)^.3 = (ab)^(3/10) = [(ab)^3]^(1/10)

OpenStudy (anonymous):

so it works for decimals

OpenStudy (anonymous):

it just doesnt work for irrationals. no proof will unless you use least upperbound property

OpenStudy (anonymous):

but how do we find (a^3 * b^3)^0.1 ?

OpenStudy (anonymous):

and show that it is equal to a^0.3 * b^0.3

OpenStudy (anonymous):

(a^3 * b^3)^0.1 = (a^3 * b^3)^(1/10) = [(a^3 * b^3)^1]^(1/10)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:):):):)

OpenStudy (anonymous):

I will borrow my professor's answer book and see the one-liner answer for myself :-D

OpenStudy (anonymous):

so we have to find fractions in general (ab)^(1/c) = a^(1/c)*b^(1/c) , prove that

OpenStudy (anonymous):

but thats easy , 1/c = c^-1

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

(a^3 * b^3)^0.1 = (a^3*b^3) ^(10^-1) = [(a^3 * b^3)^10]^-1

OpenStudy (anonymous):

mistake there.

OpenStudy (anonymous):

(a^3 * b^3)^0.1 (a^3*b^3)^1/10 = (a^3*b^3) ^(10^-1) [(a^3 * b^3)^10]^-1 [a^30 *b^30]^-1

OpenStudy (anonymous):

10^-1 is not equal to 10*-1

OpenStudy (anonymous):

1/10 = 10^-1

OpenStudy (anonymous):

true

OpenStudy (anonymous):

but that doesn't mean that a^{10^-1} is equal to (a^{10})^{-1}

OpenStudy (anonymous):

(a^3 * b^3)^0.1 (a^3*b^3)^1/10 = (a^3*b^3) ^(10^-1) [(a^3 * b^3)^10]^-1 [a^30 *b^30]^-1 1/[a^30*b^30] 1/a^30 * 1/b^30 a^-30 * b^-30

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

yes it does because you can use the rule (ab)^n = a^n*b^n

OpenStudy (anonymous):

which we proved for n>0 , for n = 0 it is trivial

OpenStudy (anonymous):

0^{-1} is undefined :-P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

we assumed that a>0, b>0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so it works, but theres a few steps

OpenStudy (anonymous):

first we show that (ab)^n = a^n*b^n, for n>0, a>0,b>0 . then we show that (ab)^(1/n) = a^(1/n)*b^(1/n)

OpenStudy (anonymous):

then we can so that (ab)^(p/q) = ...

OpenStudy (anonymous):

but still, \[a^{\frac{1}{n}} \neq ((a)^n)^{-1}\]

OpenStudy (anonymous):

i didnt write that

OpenStudy (anonymous):

you idiot, i wrote

OpenStudy (anonymous):

a^(1/n) = [ a^(n)]^-1

OpenStudy (anonymous):

you wrote this: = (a^3*b^3) ^(10^-1) [(a^3 * b^3)^10]^-1

OpenStudy (anonymous):

and it's wrong :-D

OpenStudy (anonymous):

a^(1/n) = a^ [ n^-1] =

OpenStudy (anonymous):

no thats right, youre an idiot

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

(a^3b^3)^(1/10) = (a^3*b^3) ^(10^-1)= [(a^3 * b^3)^10]^-1 get a flipping clue

OpenStudy (anonymous):

alright, use the fact that \[a^{n^{-1}} = (a^{n})^{-1} \]to find the square root of 2

OpenStudy (anonymous):

because that's what you essentially wrote

OpenStudy (anonymous):

no i didnt write that, i have a carat you left out

OpenStudy (anonymous):

(a^3*b^3) ^(10^-1)= [(a^3 * b^3)^10]^-1 I didn't leave any carets out

OpenStudy (anonymous):

one sec, brb

OpenStudy (anonymous):

mine is right, yours is wrong

OpenStudy (anonymous):

nope, you've made an error in one of your statements

OpenStudy (anonymous):

ok exponentiation is not associative, big whoop

OpenStudy (anonymous):

2^(2^3) != (2^2)^3